# Media Coverage of War Victims: Journalistic Biases in Reporting on Israel and Gaza

### v2

@@ Abstract
October 7th 2023 marked the start of a war against Gaza, which is considered one of the most devastating wars in modern history and has led to a stark attitudinal divide within and between countries. To investigate the role of media bias in reporting on this asymmetrical warfare, we analyzed over 14,000 news articles published during the first year of war in three Western (The New York Times, BBC, CNN) and one non-Western English-language outlets (Al Jazeera English). Exploring the media narratives concerning Israeli and Palestinian victims experiencing hardship, we found three systematic biases in Western media. 1) Compared to Palestinian victims, represented mainly as undifferentiated collectives, Israeli victims were more likely to be portrayed as identifiable individual human beings. 2) Despite the striking difference in all forms of hardship (casualties, displacement, etc.), Western journalists created a false balance, equating Israeli and Palestinian suffering, by persistently referring back to the 7th of October massacre, even in the absence of new events involving Israeli victims. 3) When reporting on numbers of Palestinian (vs. Israeli) victims, journalists used language that casts doubt about the credibility of the information and the reputation of the source providing it, thereby selectively undermining the reader’s trust in the information regarding Palestinian suffering. Together, our analysis reveals a series of systematic journalistic biases in high-profile Western media that are absent or greatly reduced in Al Jazeera.
@@ Introduction
The ongoing U.S.-backed war by Israel against Gaza began on October 7th, following an attack by Hamas’ military wing on Israel. The attack resulted in over 1,200 deaths, including 38 children, the taking of approximately 250 hostages, and nearly 5,500 injured [1]. The attack is considered the deadliest attack in Israel’s history. This was followed by an extensive war of Israel against Gaza that - already by the end of the first year - had produced an unprecedented humanitarian crisis [2], the destruction of over half of Gaza’s homes [3], the devastation of the school, university, and health systems, extreme deprivation, and the forcible displacement of almost the entire population [1]. After one year of war, the death toll among Palestinians in Gaza amounted to over 42,000 people [4], the majority of whom women and children, although the true death toll may exceed 186,000 once indirect deaths are taken into account [5]. Over 97,000 people were injured [4], one fourth of whom with life-changing injuries such as amputations [6]. Different from other conflicts (such as Ukraine), the population is trapped in Gaza due to the 17-year-old air, sea and land blockade. The war has been particularly dire on children. In the first twelve months alone, over 14,000 had been killed [7], at least 17,000 were unaccompanied [8], and 90% suffered acute malnutrition [9]. According to Save the Children, more than 10 children each day had one or both legs amputated [10], and by the time of the writing, the humanitarian crisis has tragically deteriorated and multiple organizations, including the United Nations [11], Human Rights Watch [12], and Amnesty International [13], have declared that a genocide is taking place.
@@
Against this backdrop, the present study examines how such a humanitarian disaster and information bottleneck are reflected in international media coverage. The current analysis, informed by psychological research and theory, focuses on subtle media biases in war journalism. Specifically, we examine three critical facets: (1) (De-)humanizing reporting, assessing to what extent victims are portrayed as individual human beings or as undifferentiated collectives. Thus, we look at the balance between group-level reports and individualized stories, which are more likely to evoke empathy. (2) Quantifying the human costs of war, exploring whether media reports present numbers in a meaningful and proportional manner, accurately reflecting the distribution of victims (and casualties more specifically) over the 12-month period examined. (3) Reporting on child victims, focusing on how children are individualized and featured in media coverage, as well as how their suffering is quantified.
@@ Results
In this study, we analyze over 14,000 news articles published between October 7, 2023, and October 7, 2024, from The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE). We focus solely on the textual content, deliberately excluding other media formats like videos, podcasts and transcripts to maintain consistency in the analysis. For a detailed breakdown of the data composition and collection methodology, refer to the Materials and Methods’ Data subsection. As we were interested in subtle biases, we purposefully focused on highly professional news outlets, while excluding tabloids or other popular news sources where blatant partisanship could be expected.
@@
With this data, we conducted our key analyses. We investigated how media outlets handled (de-)humanizing reporting, quantified the human costs of war, and reported on child victims. In order to establish bias, we compared media reporting, wherever possible, with objective standards such as actual casualty rates, as explained in greater detail below. This allowed us to uncover nuanced media biases across different sources. Importantly, with the term ”bias”, we are not assuming any form of intentionality or malevolence, but describing observable asymmetries in coverage that reflect underlying reporting tendencies. The possible reasons for the occurrence of such biases will be discussed in later sections.
(De-)humanizing the Victims: Individualized vs Category-based Reporting
War journalism has been criticized for its primary focus on frontline reporting, often neglecting the human consequences for the civilian population, and for children in particular [20]. The media may either quantitatively give little space to civilian victims or report on them in a way that conceals their full human qualities (dehumanization), including their intellect, their culture, their morality, their feelings, and their individuality. One way in which journalists may downplay the civilian experience of war is by reporting group-level experiences rather than individual stories. Social scientists such as Slovic [21] and Schelling [22] have long observed that people often remain indifferent to mass atrocities such as war, mass murder, and genocide, yet are profoundly moved by individual stories. This psychological tendency is captured by what scholars call identifiable victim effect [23]: people feel more compassion toward single victims and are more likely to take action (for instance, through donations) when text or images describe the suffering of a single individual rather than that of larger groups. Compassion quickly fades as the number of victims increases [21], with the exception of very cohesive and interconnected small groups such as families [24]. For instance, in a classical study by Kogut and Ritov [25], a single child in need of a costly life-saving treatment elicited greater distress and received greater donations than a group of sick children with identical needs. This phenomenon is driven by different mechanisms, most notably that identifiable victims attract greater attention and evoke greater emotional reactions (such as sympathy or distress) than statistical victims or large groups [21]. Also, by providing personalized information (such as age or name), as is typically done for single identifiable victims, narratives tend to be more vivid, which in turn induces a sense of familiarity and closeness. Unsurprisingly, the identifiable victim effect is more powerful when the victim is not responsible for their plight [23]. This is particularly relevant when reporting on children who, by definition, cannot be held responsible for an armed conflict or for the causes that led up to it. As such, by examining both the degree of individualization in war reporting and the specific attention devoted to children, our analysis seeks to uncover whether media narratives amplify or diminish the humanity of those most powerless in the face of war.
@@
We began our analysis by identifying all instances in which individuals or groups were mentioned in news articles across our four media sources. To do this, we employed a large language model (LLM) to extract every reference to civilians, whether individual or collective, and classify each instance as pertaining to either the Palestinian or Israeli side. Specifically, we instructed the model to identify individualized civilian stories that highlight personal experiences of hardship, such as: “Noralin ‘Nataly’ Babadilla, 60, was visiting Kibbutz Nirim, to celebrate the community’s 70th anniversary. Her husband was murdered….” [26]. These types of narratives humanize the conflict by centering on a specific, emotionally resonant character. In addition, we asked the model to categorize the most severe type of hardship being described, enabling us to analyze not just whether individuals are mentioned, but also how their suffering is framed. Group-level references, which are typically broader and less personalized, included generalized hardships, such as “With many Israelis traumatized by the attack…” [27], statistical accounts, such as “More than 40,738 people have been killed in Gaza” [28], and event descriptions, like “Gazans are dying of starvation” [29]. A detailed explanation of how these classifications were extracted and validated is provided in Supplementary Note 1.
@@
Next, we examined the extent to which each side is represented through individualized versus group narratives. As shown in Figure 2a, Western media outlets, BBC, CNN, and NYT, consistently featured higher ratios of individualized to group mentions for the Israeli compared to the Palestinian side. This asymmetry was most pronounced in NYT, where Israeli individuals were referenced once for every six group mentions (ratio = 0.16), whereas Palestinians were individualized only once for every twelve group mentions (ratio = 0.08). Similar patterns emerged for BBC (0.20 vs. 0.12) and CNN (0.13 vs. 0.09), though the disparity was less stark (see Supplementary Table 3). In contrast, AJE showed a more balanced pattern, with nearly equal ratios for both sides (0.09 for Palestinians vs. 0.07 for Israelis). This suggests that Israeli individuals were more frequently singled out in Western media coverage, while Palestinians were more often portrayed as part of an undifferentiated collective. We formalized this asymmetry using a Generalized Estimating Equations (GEE) regression analysis (see Materials and Methods, subsection “Testing Individualization Bias with GEE Models”), which confirmed that Western outlets were significantly more likely to individualize the Israeli side in their reporting. Specifically, we find that the odds of a reported story being individualized are 13.6% lower for Palestinian instances compared to Israeli ones in the case of BBC, 27.0% lower in the case of CNN, and 38.9% lower in the case of NYT. No statistically significant difference was observed for AJE.
Imbalance in Casualty Reporting
@@
When examining the ratio of Palestinian to Israeli casualty-related stories, as shown on the left side of Figure 2b, a stark contrast was revealed between AJE and Western outlets. AJE published roughly eight Palestinian casualty stories for every Israeli story, while BBC, CNN, and NYT maintained a close to one-to-one ratio. NYT even showed a slight tilt in the opposite direction, reporting eight Palestinian stories for every ten Israeli ones. At first glance, this might suggest a pro-Palestinian bias in AJE’s reporting. However, when we compared these ratios to the actual casualty figures, used here as a crude baseline for contextualizing media attention, a more complex picture emerged. As shown on the right side of Figure 2b, for every Israeli casualty, there were roughly 20 Palestinian casualties by the end of the first year of war. This stark disparity in actual casualties reflects the marked power differential between the two groups, and contrasts with the seemingly “balanced” reporting in Western media. In theory, if all events were covered equally and without selection bias, one would expect a proportional discrepancy in the number of stories reflecting the underlying reality. Instead, the reporting style of Western media and, to a lesser degree of AJE, resembles what is known as “false balance”, namely the tendency to create the impression that two sides are equivalent when, in reality, they are not (as will be explained in more detail in the discussion section).
@@
It is important to note, however, that one would not necessarily expect a perfect match between actual casualty rates and casualty reporting, indeed, such correspondence was absent in all outlets, including AJE. Some degree of over-reporting of the events of October 7th compared to the following twelve months would be expected on the basis of well-documented psychological mechanisms. First, the October 7th massacre was extremely salient, involving a large number of victims in a short time span. Research on the availability heuristic shows that a single-mass casualty event tends to attract disproportionate attention and to be remembered more vividly than a higher, but gradually accumulating, death tolls [30]. Second, events that occur first in time or space tend to receive greater attention than those that follow (e.g., the ballot-order effect [31]. Thus, to the extent that the October 7 attack is perceived as the “starting point” of the conflict, it is understandable that it received over-proportional coverage. Nevertheless, such cognitive and temporal biases cannot account for the extent to which reporting equated casualties on both sides. For instance, while the 38 Israeli children killed on October 7th might be expected to draw more attention than the 38 Palestinian children killed, on average, each day over the subsequent year, it would be implausible for the cumulative 14,000 Palestinian child deaths to receive equivalent attention as those 38 Israeli deaths.
Sustaining Parity in the Face of Discrepancy
@@
This led us to a critical line of inquiry: How do Western media outlets sustain parity in reporting casualty-related stories between both sides when the actual number of casualties is so drastically imbalanced? To investigate this, we analyzed the content of Israeli casualty-related stories published by Western media (generated using Prompt 2 in Supplementary Note 1). Strikingly, we found that 92% of these stories were related to the events of October 7, as shown in Figure 2c. This concentration persisted even in stories published several months after the attack. In other words, a single moment in time, the initial event, was repeatedly covered to sustain a consistent volume of Israeli casualty narratives, thereby flattening the temporal and numerical asymmetry in actual deaths.
@@
A more detailed explanation of our date-selection procedure and findings, including additional quotes, is provided in Supplementary Note 2. The number of Israeli stories, related to October 7, published by each outlet on the dates examined can be found in Supplementary Table 4.
Quantifying Empathy-Related Narrative Elements
To assess whether the qualitative imbalance we observed persisted across the full year of coverage, we quantified differences in narrative elements that can influence reader empathy, using the validated Human Empathy and Narrative Taxonomy (HEART) framework [50]. Specifically, we applied two HEART metrics: (i) Vividness of Emotions and (ii) Plot Volume (the main predictors of perceived empathy in the HEART model) to all hardship-related individualized stories in our corpus. While these metrics are not direct measures of empathy, prior research shows they are key stylistic features that contribute to how much empathy a narrative may elicit. Scores for each metric were first calculated at the story level and then aggregated by side within each article, given that articles often addressed both sides. For each metric, we then calculated a difference score by subtracting the Israeli aggregate score from the Palestinian one, with positive values indicating that these empathy-related narrative elements were more prominent in Palestinian stories, and negative values indicating the reverse. For full details on the scoring procedure, see “Measuring Narrative Elements Associated with Empathy in News Texts” in the Materials and Methods.
@@
The Plot Volume results show the same pattern, including the two spikes, with Israeli-focused stories in those periods receiving even higher relative scores than in the Vividness of Emotions metric. At the same time, it is worth noting that, on average, stories about Palestinians were written with greater emotional vividness across outlets (albeit less so for NYT), which is not surprising given the scale of Palestinian suffering. What is striking, however, is that even outlets such as BBC and CNN, while often writing about Palestinians in a more emotional way, still devoted disproportionate attention to Israeli narratives during key hostage-related events. Taken together, these findings suggest that the asymmetry lies less in the tone itself than in the balance of coverage: Western outlets often create a false equivalence by giving equal or greater narrative weight to Israeli experiences, despite the vastly different realities of suffering on the ground.
False Balance in the Coverage of Child Victims
@@
Given the high vulnerability of children and the fact that they cannot be held responsible for armed conflict nor for the election of leaders who promote armed conflicts, it is not surprising that children have a special status in international humanitarian law (among which the UN Convention on the Rights of the Child [51]). Similarly, lay people value the life of the young higher than that of adults [52], feel a greater moral obligation to help children [53], and are more likely to assist them [54, 55]. Children (including older children) are prioritized in helping decisions because they evoke greater nurturing concern and compassion [56], because they are generally not held responsible for their fate, and because they have, on average, a greater number of years ahead, which makes helping them a rational choice. As such, it is essential to understand how much attention the different media sources pay to child victims of war. We therefore extended our analyses of individualized reporting to specifically track mentions of children.
@@
On average, the four media outlets dedicated roughly 46.5% of their individualized stories to children (AJE 47%, BBC 45%, CNN 50%, NYT 44%; Supplementary Table 5a). Taken at face value, this looks like responsible reporting: the media gave children more prominence than their actual share of casualties (about 32%), which seems appropriate given children’s heightened vulnerability and moral salience. However, when the data are broken down by side, the picture changes. Palestinian children do take up a commendable share of stories—roughly in line with their 34% share of deaths—but Israeli children receive far more attention than their casualty numbers would suggest. Although they make up only 3% of Israeli deaths, their stories account for as much as 44% of child-related individualized reporting in CNN, with similarly elevated shares in BBC (36%) and NYT (40%). This disproportionate emphasis mirrors the false balance documented in our earlier analyses, where the relative weight given to Israeli and Palestinian suffering diverged sharply from the underlying reality.

### v3

@@ Abstract
October 7, 2023 marked the start of a war against Gaza, one of the most devastating conflicts in modern history, which quickly produced a stark global attitudinal divide. To examine the role of media bias in shaping public understanding of this asymmetrical war, we analyzed more than 14,000 news articles published during its first year across three major Western outlets (The New York Times, BBC, CNN) and one non-Western English-language outlet (Al Jazeera English). Focusing on media narratives surrounding Israeli and Palestinian victims, we identify three systematic biases in Western coverage: (1) Identifiable Victim Reporting: Israeli victims were substantially more likely to be depicted as identifiable individuals, whereas Palestinian victims were predominantly represented as undifferentiated collectives. (2) Equalization Bias: Despite the profound asymmetry in casualties, displacement, and other forms of suffering, Western reporting repeatedly invoked the October 7 attacks to equalize Israeli and Palestinian victimhood, even in the absence of new Israeli-casualty events. (3) One-sided Doubt Casting: Journalists disproportionately used language that casts doubt on the credibility of casualty figures and the reliability of sources when reporting Palestinian (vs. Israeli) victim counts, selectively undermining trust in information about Palestinian suffering. Across all three phenomena, these patterns were either absent or greatly attenuated in Al Jazeera English. Taken together, our analysis uncovers a coherent set of systematic biases in high-profile Western media coverage of the Gaza war, with implications for how global audiences come to understand and morally evaluate the conflict.
@@ Introduction
The ongoing U.S.-backed war by Israel against Gaza began on October 7th, following an attack by Hamas’ military wing on Israel. The attack resulted in over 1,200 deaths, including 38 children, the taking of approximately 250 hostages, and nearly 5,500 injured [1]. The attack is considered the deadliest attack in Israel’s history. This was followed by an extensive war of Israel against Gaza that - already by the end of the first year - had produced an unprecedented humanitarian crisis [2], the destruction of over half of Gaza’s homes [3], the devastation of the school, university, and health systems, extreme deprivation, and the forcible displacement of almost the entire population [1]. After one year of war, the death toll among Palestinians in Gaza amounted to over 42,000 people [4], the majority of whom women and children. However, independent estimates suggest that the true death toll may greatly exceed these official statistics, both for violent mortality and when non-violent excess death rates are taken into account [5, 6]
Over 97,000 people were injured [4], one fourth of whom with life-changing injuries such as amputations [7]. Different from other conflicts (such as Ukraine), the population was trapped in Gaza due to the 17-year-old air, sea and land blockade. The war has been particularly dire on children. In the first twelve months alone, over 14,000 had been killed [8], at least 17,000 were unaccompanied [9], and 90% suffered acute malnutrition [10]. According to Save the Children, more than 10 children each day had one or both legs amputated [11], and by the time of the writing, the humanitarian crisis has tragically deteriorated and multiple organizations, including the United Nations [12], Human Rights Watch [13], and Amnesty International [14], have declared that a genocide is taking place.
@@
Against this backdrop, the present study examines how such a humanitarian disaster and information bottleneck are reflected in international media coverage. The current analysis, informed by psychological research and theory, focuses on subtle media biases in war journalism. Specifically, we examine three potential journalistic biases that may have played a role in creating or sustaining attitudinal divides:
  1. 1.
    Identifiable victim reporting: War journalism has been criticized for its primary focus on frontline reporting, often neglecting the human consequences for the civilian population, and for children in particular [21]. The media may either quantitatively give little space to civilian victims or report on them in a way that conceals their full human qualities (dehumanization), including their intellect, their culture, their morality, their feelings, and their individuality. One way in which journalists may downplay the civilian experience of war is by reporting group-level experiences rather than individual stories. Social scientists such as Slovic [22] and Schelling [23] have long observed that people often remain indifferent to mass atrocities such as war, mass murder, and genocide, yet are profoundly moved by individual stories. This psychological tendency is captured by what scholars call identifiable victim effect [24]: people feel more compassion toward single victims and are more likely to take action (for instance, through donations) when text or images describe the suffering of a single individual rather than that of larger groups. Compassion quickly fades as the number of victims increases [22], with the exception of very cohesive and interconnected small groups such as families [25]. For instance, in a classical study by Kogut and Ritov [26], a single child in need of a costly life-saving treatment elicited greater distress and received greater donations than a group of sick children with identical needs. This phenomenon, mostly investigated in experimental work, is driven by different mechanisms, most notably that identifiable victims attract greater attention and evoke greater emotional reactions (such as sympathy or distress) than statistical victims or large groups [22]. Also, by providing personalized information (such as age or name), as is typically done for single identifiable victims, narratives tend to be more vivid, which in turn induces a sense of familiarity and closeness. To examine the relative humanization of victims through individual stories, we therefore analyzed the degree to which Israeli and Palestinian civilian victims were portrayed through individualized (rather than collective) narratives. In the absence of bias, one would expect that the ratio of individualized to group-based narratives would be approximately the same for both sides. To the best of our knowledge, this is the first large-scale media analysis of individualizing vs. group-level reporting in the context of war (for a small-scale qualitative analysis on identifiable victim reporting, see [27]).
  2. 2.
    Equalizing bias: The second journalistic bias investigated here refers to the tendency to deny hierarchies in human suffering. In asymmetrical conflict situations, in which victims on one side greatly outnumber those on the other, one might expect media coverage to roughly mirror this difference, whereas stark deviations from this baseline may suggest that not all lives are equally news-worthy. We coined the term Equalization Bias to describe the tendency to dedicate equal media attention or space to two groups that, in reality, show very different levels of human suffering. For instance, giving the same coverage to Bosnian and Serbian victims in the Yugoslav war would constitute a case of equalizing bias. The equalization bias resembles, and was inspired by, the concept of false balance [28], an established journalistic bias, where two positions are presented as epistemically equivalent and receive equal space or attention (e.g. airtime) despite differences in actual validity and trustworthiness, such as climate skeptics and climate scientists or pro- and anti-vaccination proponents [28, 29, 30]. Both, the false balance and the equalization bias create a false equivalence between non-comparable objects, but whereas false balance revolves around the question of validity of information, the equalization bias proposed here alludes to the allocation of narrative resources to different groups (e.g., number of stories are told, space devoted, etc) that is not proportionate of the actual scale of human suffering. Furthermore, given the novelty of the proposed bias, we complement our large-scale media analysis with a controlled experiment designed to identify the psychological mechanism that may give rise to this bias.
  3. 3.
    One-sided doubt-casting: The third bias refers to linguistic devices (including hedges such as “allegedly”) that signal that a reported fact is unconfirmed and, therefore, that the claim is uncertain (e.g., Bailey et al., 2014 [31]). Such statements imply a lack of verifiability or consensus and, hence, are likely to induce doubt in the reader’s mind. In the context of war or natural disasters, where actual victim numbers can not easily be established or independently verified, journalists necessarily rely on external sources, typically governmental agencies such as the police or the health ministries of the countries involved. Whereas doubt-casting language is generally considered a sign of journalistic integrity and may reflect high ethical standards in situations of uncertainty, it becomes problematic when applied in a one-sided manner. To uncover potential bias, we therefore investigate the degree to which doubt-casting is used to either undermine the credibility of the source of information (e.g., according to the Hamas-controlled health ministry) or to question the reliability of the victim numbers reported in an article (such as allegedly; reportedly; appear).
Having outlined the three forms of journalistic bias, we next clarify how they are explored in the present study. Importantly, throughout this work we use the term “bias” to describe systematic asymmetries or deviations in media coverage relative to objective baselines, such as actual casualty rates. The concept is applied to all three phenomena discussed above, without implying any form of intentional distortion or malevolent intent on the part of journalists. Possible explanations for the emergence of such patterns are discussed in later sections.
Our analyses focus on the human consequences of war for the civilian population (including children), that constitute the overwhelming majority of victims on both, the October 7 attack (69% of all victims) and in the subsequent war on Gaza (estimated to be above 80% among Palestinian victims [32]). To investigate the three proposed biases, we employ two complementary analytical approaches: one centered on narrative representation and the other on quantitative reporting.
  • Narrative representation examines how victims are portrayed in media stories, whether as individualized human beings or as undifferentiated collectives, capturing degrees of (de-)humanization in reporting.
  • Quantitative reporting investigates how numbers are used to convey the human costs of war, assessing whether the distribution of attention and reported casualties reflects the actual scale, the asymmetry of suffering across groups and whether such numbers are questioned.
Together, these two analytical dimensions allow us to systematically assess how subtle reporting tendencies may shape perceptions of conflict and contribute to the broader attitudinal divides observed across the international media landscape.
@@ Results
In this study, we analyze over 14,000 news articles published between October 7, 2023 and October 7, 2024 by four major international media outlets: The New York Times (NYT), BBC, CNN, and Al Jazeera English (AJE). To ensure analytical consistency and cross-outlet comparability, we restrict our corpus to textual news content and exclude videos, podcasts, and transcripts. A detailed description of the outlet selection criteria, temporal scope, and data collection procedures is provided in the Materials and Methods (Data Selection and Sample Construction subsection).
@@
Using this dataset, we examined the three biases outlined in the Introduction. We began with narrative representation, focusing on the first two biases, identifiable victim reporting and equalizing bias, by asking whose stories are individualized, whose remain collective, and how casualty-related reporting reflects underlying asymmetries in victimization.
Narrative Representation as an Indicator of Identifiable Victim and Equalizing Bias
Narrative representation concerns how civilian victims are made visible (or remain obscured) in news coverage. Individualized stories tend to humanize victims and elicit empathy, whereas collective or statistical references do so to a much lesser extent. Examining these narrative choices enables us to identify two forms of bias: whether some victims are more likely to be individualized (identifiable victim bias) and whether narrative space is distributed in ways that obscure large disparities in suffering (equalizing bias).
Identifiable Victim Bias: Individualized vs Collective Reporting
@@
We began by examining the identifiable victim bias, that is, the tendency to portray some victims as individualized persons while others are presented as undifferentiated collectives. To do this, we employed a large language model (LLM) to extract every reference to civilians, whether individual or collective, and classify each instance as pertaining to either the Palestinian or Israeli side. Specifically, we instructed the model to identify individualized civilian stories that highlight personal experiences of hardship, such as: “Noralin ‘Nataly’ Babadilla, 60, was visiting Kibbutz Nirim, to celebrate the community’s 70th anniversary. Her husband was murdered….” [33]. These types of narratives humanize the conflict by centering on a specific, emotionally resonant character. In addition, we asked the model to categorize the most severe type of hardship being described, enabling us to analyze not just whether individuals are mentioned, but also how their suffering is framed. Group-level references, which are typically broader and less personalized, included generalized hardships, such as “With many Israelis traumatized by the attack…” [34], statistical accounts, such as “More than 40,738 people have been killed in Gaza” [35], and event descriptions, like “Gazans are dying of starvation” [36]. A detailed explanation of how these classifications were extracted and validated is provided in Supplementary Note 1.
@@
Next, we examined the extent to which each side is represented through individualized versus group narratives. As shown in Figure 2a, Western media outlets, BBC, CNN, and NYT, consistently featured higher ratios of individualized to group mentions for the Israeli compared to the Palestinian side. This asymmetry was most pronounced in NYT, where Israeli individuals were referenced once for every six group mentions (ratio = 0.16), whereas Palestinians were individualized only once for every twelve group mentions (ratio = 0.08). Similar patterns emerged for BBC (0.20 vs. 0.12) and CNN (0.13 vs. 0.09), though the disparity was less stark (see Supplementary Table 3). In contrast, AJE showed a more balanced pattern, with similar ratios for both sides (0.09 for Palestinians vs. 0.07 for Israelis). This suggests that Israeli individuals were more frequently singled out in Western media coverage, while Palestinians were more often portrayed as part of an undifferentiated collective.
We formalized this asymmetry using a Generalized Estimating Equations (GEE) regression analysis (see Materials and Methods, subsection “Testing Individualization Bias with GEE Models”), which confirmed that Western outlets were significantly more likely to individualize the Israeli side in their reporting. Specifically, we find that the odds of a reported story being individualized are 14.7% lower for Palestinian instances compared to Israeli ones in the case of BBC, 28.7% lower in the case of CNN, and 41.3% lower in the case of NYT. No statistically significant difference was observed for AJE. Taken together, these results provide clear evidence of an identifiable victim bias in Western coverage, whereby Israeli civilians are more often presented as fully individuated persons, whereas Palestinians are more frequently portrayed as part of an anonymous collective.
Equalizing Bias in Casualty Reporting
@@
After identifying casualty-related stories, we next assessed the presence of equalizing bias by comparing how frequently outlets reported on Palestinian versus Israeli casualties. As shown on the left side of Figure 2b, a stark contrast was revealed between AJE and Western outlets. AJE published roughly eight Palestinian casualty stories for every Israeli story, while BBC, CNN, and NYT maintained a close to one-to-one ratio. NYT even showed a slight tilt in the opposite direction, reporting eight Palestinian stories for every ten Israeli ones. At first glance, this might suggest a pro-Palestinian bias in AJE’s reporting. However, when we compared these ratios to the actual casualty figures, used here as a crude baseline for contextualizing media attention, a more complex picture emerged. As shown on the right side of Figure 2b, for every Israeli casualty, there were roughly 20 Palestinian casualties by the end of the first year of war. This stark disparity in actual casualties reflects the marked power differential between the two groups, and contrasts with the seemingly “balanced” reporting in Western media. In theory, if all events were covered equally and without selection bias, one would expect a roughly proportional discrepancy in the number of stories reflecting the underlying reality. Instead, the reporting style of Western media and, to a lesser degree of AJE, is coherent with our phenomenon of equalizing bias, namely the tendency to create the impression that two sides are equivalent when, in reality, they are not. In particular, Western outlets create a false equivalence by giving equal narrative weight to Israeli experiences, despite the vastly different realities of suffering on the ground.
@@
It is important to note, however, that one would not necessarily expect casualty reporting to mirror casualty rates perfectly. In fact, no outlet, including AJE, displayed a fully proportional relationship. Some degree of over-reporting of the events of October 7th relative to the subsequent months is psychologically plausible. First, the October 7th massacre was extremely salient, involving a large number of victims in a short time span. Research on the availability heuristic shows that single-mass casualty events tend to attract disproportionate attention and are remembered more vividly than larger, but gradually accumulating, death tolls [37]. Second, events that occur first in time or space often receive greater attention than those that follow; a pattern observed in phenomena such as the ballot-order effect [38]. Thus, if October 7 is perceived as the “starting point” of the war, some degree of disproportionate coverage would be understandable. Nevertheless, such cognitive and temporal biases cannot account for the extent to which media reporting equated casualties on both sides. For instance, while the 38 Israeli children killed on October 7th might be expected to draw more attention than 38 Palestinian children killed, on average, each day over the subsequent year, it would be implausible for the cumulative 14,000 Palestinian child deaths over the course of the year to warrant narrative attention equivalent to those 38 Israeli deaths. This pattern is consistent with equalizing bias: a structural tendency to allocate roughly equal narrative space to two sides despite profound asymmetries in underlying human suffering. Such patterns may be shaped, in part, by asymmetries in identification with conflict parties, which we next attempted to test experimentally.
Experimental Evidence on Identification as a Mechanism of Equalizing Bias
To examine whether equalizing bias can arise from social identification alone, i.e. the degree to which individuals perceive a group as part of their social self and feel attached to it, we conducted a controlled experiment in which participants in the US and UK were presented with a fictitious asymmetric geopolitical conflict between two countries differing in military strength and civilian death tolls (10 vs. 40). Participants were randomly assigned to adopt the role of a journalist aligned with the stronger country, the weaker country, or an independent international outlet (see Supplementary Note 2 for full details). Participants reviewed brief profiles of 50 civilian casualties (10 from the stronger country; 40 from the weaker country) and were asked to select 10 victims whose stories they would feature in their news articles. In the absence of bias (random selection), proportional coverage would yield, on average, two strong-country victims and eight weak-country victims.
Consistent with the equalizing bias observed in our media analysis, participants assigned to the stronger-country condition substantially over-selected strong-country victims (mean of 3.7 rather than 2) relative to the proportional benchmark. Participants assigned to the weaker-country condition did not show such over-selection (mean of 1.8 rather than 2), while independent participants occupied an intermediate position.
Critically, identification predicted selection behavior: the more participants identified with the strong country, the more strong-country victims they selected ($r=.43$, $p<.001$). Further supporting the critical role of identification, a mediation model (see Supplementary Note 2) demonstrated that the equalizing bias is driven by journalists’ identification with the country to which they were assigned. Together, these findings provide causal evidence that equalizing bias can arise from asymmetries in social identification, lending support to the interpretation that the patterns observed in BBC, CNN, and NYT coverage are shaped by stronger identification with Israel.
Sustaining Equalizing Bias in the Face of Discrepancy
@@
This led us to a critical line of inquiry: How do Western media outlets sustain parity in reporting casualty-related stories between both sides when the actual number of casualties is so drastically imbalanced? To investigate this, we analyzed the content of Israeli casualty-related stories published by Western media (generated using Prompt 2 in Supplementary Note 1). Strikingly, we found that 92% of these stories were related to the events of October 7, as shown in Figure 2c. This concentration persisted even in stories published several months after the attack. In other words, a single moment in time, the October 7 event, was repeatedly revisited to maintain a steady volume of Israeli casualty narratives, thereby flattening the temporal and numerical asymmetry in actual deaths to sustain equalizing bias.
@@
A more detailed explanation of our date-selection procedure and findings, including additional quotes, is provided in Supplementary Note 3. The number of Israeli stories, related to October 7, published by each outlet on the dates examined can be found in Supplementary Table 4.
Empathy-Related Narrative Elements in Individualized Stories
Given that Israeli victims are more likely to be portrayed as identifiable individuals than Palestinian victims, the question arises whether the individualized stories about the former are also told in a more empathic narrative style. To explore this, we quantified differences in narrative elements that can influence reader empathy, using the validated Human Empathy and Narrative Taxonomy (HEART) framework [57]. Specifically, we applied two HEART metrics: (i) Vividness of Emotions and (ii) Plot Volume (the main predictors of perceived empathy in the HEART model) to all hardship-related individualized stories in our corpus. While these metrics are not direct measures of empathy, prior research shows they are key stylistic features that contribute to how much empathy a narrative may elicit. Scores for each metric were first calculated at the story level and then aggregated by side within each article, given that articles often addressed both sides. For each metric, we then calculated a difference score by subtracting the Israeli aggregate score from the Palestinian one, with positive values indicating that these empathy-related narrative elements were more prominent in Palestinian stories, and negative values indicating the reverse. For full details on the scoring procedure, see “Measuring Narrative Elements Associated with Empathy in News Texts” in the Materials and Methods.
@@
The Plot Volume results show the same pattern, including the two spikes, with Israeli-focused stories in those periods receiving even higher relative scores than in the Vividness of Emotions metric. At the same time, it is worth noting that, on average, stories about Palestinians were written with greater emotional vividness across outlets (albeit less so for NYT), which is not surprising given the scale of Palestinian suffering. Taken together, the above findings suggest that the asymmetry lies less in the narrative tone than in the relative frequency with which individualized stories are told in the first place: In Western media, Israeli suffering is more likely to be told through individual stories, whereas collective statements abound for Palestinian victims. Yet, once a journalist decides to report individualized stories about Palestinians, these stories are, on average, no less empathic towards the victims than those told about Israelis.
Children as a test case of Equalizing Bias
@@
Children offer a particularly revealing test case for equalizing bias. Given the high vulnerability of children and the fact that they cannot be held responsible for armed conflict nor for the election of leaders who promote armed conflicts, it is not surprising that children have a special status in international humanitarian law (among which the UN Convention on the Rights of the Child [58]). Similarly, lay people value the life of the young higher than that of adults [59], feel a greater moral obligation to help children [60], and are more likely to assist them [61, 62]. Children (including older children) are prioritized in helping decisions because they evoke greater nurturing concern and compassion [63], because they are generally not held responsible for their fate, and because they have, on average, a greater number of years ahead, which makes helping them a rational choice. As such, one might expect coverage to track the underlying disparity in child casualties especially closely. We therefore extended our analyses of individualized reporting to specifically track mentions of children.
@@
On average, the four media outlets dedicated roughly 46.5% of their individualized stories to children (AJE 47%, BBC 45%, CNN 50%, NYT 44%; Supplementary Table 5a). Taken at face value, this looks like responsible reporting: the media gave children more prominence than their actual share of casualties (about 32%), which seems appropriate given children’s heightened vulnerability and moral salience. However, when the data are broken down by side, the picture changes. Palestinian children do take up a commendable share of stories—roughly in line with their 34% share of deaths—but Israeli children receive far more attention than their casualty numbers would suggest. Although they make up only 3% of Israeli deaths, their stories account for as much as 44% of child-related individualized reporting in CNN, with similarly elevated shares in BBC (36%) and NYT (40%). This disproportionate emphasis mirrors the equalization bias documented in our earlier analyses, where the relative weight given to Israeli and Palestinian suffering diverged sharply from the underlying reality.